# Algorithm Design and Prototype Validation for Reconfigurable Intelligent Sensing Surface: Forward-Only Transmission

### v2

@@ V-C Sensing-assisted Communication for Interference Suppression
In this section, we further evaluate the performance of interference suppression [^13] For detailed implementation videos, please refer to https://drive.google.com/file/d/11SJT2nxz9Zuo5t7ZtBZR4MCcfASvaUpO/view?usp=drive_link.. The directions of the target signal and interference signal are set to $\phi^{\text{azi}}_{1}=-20^{\circ}$ and $\phi^{\text{azi}}_{2}=10^{\circ}$, respectively. As shown in Fig. 13, the interference signal is located in the side-lobe direction, leading to strong interference at the receiver. Fig. 14 presents the communication constellation diagram with interference suppression using the near-field model, compared to the counterpart scheme that only aligns with the target signal. It can be observed that the interference signal reduces the SINR and causes the constellation diagram to become more dispersed. Furthermore, the constellation diagram exhibits a pattern similar to that of non-orthogonal multiple access (NOMA), due to the power disparity between Tx. 1 and Tx. 2. However, the interference suppression scheme significantly improves the SINR, restoring the compactness of the constellation diagram. The theoretical and experimental beampatterns of both schemes are shown in Fig. 15. It is noteworthy that both the theoretical and experimental results presented in Fig. 15a and Fig. 15b demonstrate a consistent trend in interference suppression and target signal enhancement achieved through the 2-bit RISS. Specifically, a gain is observed in the direction of the target signal, while deep nulls are formed in the directions of interference. However, due to practical system imperfections such as phase shift deviations in the reflection elements and environmental disturbances, the actual level of interference suppression exhibits a noticeable deviation from the theoretical prediction. Additionally, it is also observed that the interference suppression scheme results in a slight reduction in the main-lobe gain compared to the only align scheme.
@@
(b) Robust interference suppression for interference signal, $\Delta_{\text{EVM}}=6.5897\%$.
@@
Figure 15: The beampattern of align scheme and interference suppression scheme in the near-field model. The target signal is located in $\phi_{1}^{\text{azi}}=-20^{\circ}$, while the interference signal is located in $\phi_{2}^{\text{azi}}=10^{\circ}$.
In contrast, as illustrated in Fig. 16 and Fig. 17, the performance of target enhancement and interference suppression in the far-field model is weak. To better understand the underlying cause of this phenomenon, we present theoretical 2D heatmaps under both the NF and FF schemes. It can be observed from Fig. 18 that both schemes place the Tx.2 within the main-lobe while forming nulls in the directions of Tx.1. However, the NF scheme exhibits fewer sidelobes and does not generate significant signal gain in the near region. In contrast, the FF scheme produces more pronounced sidelobes and exhibits gain along the entire path at $\phi_{2}^{azi}=10^{\circ}$. In addition, there is a slight offset between the main-lobe direction and the actual position of the Tx.2. This misalignment is caused by the combined effects of the interference suppression constraint and the 2-bit RISS quantization. Furthermore, hardware imperfections, such as phase adjustment errors in the RISS, may also be contributing factors to the inferior performance of the FF scheme compared to the NF scheme.
@@
(b) Robust interference suppression for interference signal, $\Delta_{\text{EVM}}=20.0157\%$.
@@
Figure 17: The beampattern of align scheme and interference suppression scheme in the far-field model. The target signal is located in $\phi_{1}^{\text{azi}}=-20^{\circ}$, while the interference signal is located in $\phi_{2}^{\text{azi}}=10^{\circ}$.
(a) Theoretical heatmap of NF scheme.
(b) Theoretical heatmap of FF scheme.
Figure 18: Heatmaps of both NF and FF schemes. The RISS is located at (0,0) and target signal is located in $\phi_{1}^{\text{azi}}=-20^{\circ}$, while the interference signal is located in $\phi_{2}^{\text{azi}}=10^{\circ}$.
Finally, to further investigate interference suppression, we deactivate Tx. 2, allowing the receiver to capture only the interference signal. The constellation measurements for both the align scheme and the interference suppression scheme using the near-field model are displayed in 19. With only Tx. 1 active, the constellation in 19a shows that the align scheme retains significant interference signal power. In contrast, the interference suppression scheme shows effective interference mitigation, resulting in a highly dispersed constellation diagram of the interference signal, further demonstrating the robustness of the interference suppression technique.
@@ VI Conclusion
Appendix A Hardware Information Supplement
@@
Array Size 10$\times$10 units

### v6

@@ V-C Sensing-assisted Communication for Interference Suppression
In this section, we further evaluate the performance of interference suppression [^14] For detailed implementation videos, please refer to https://drive.google.com/file/d/11SJT2nxz9Zuo5t7ZtBZR4MCcfASvaUpO/view?usp=drive_link.. The directions of the target signal and interference signal are set to $\phi^{\text{azi}}_{2}=10^{\circ}$ and $\phi^{\text{azi}}_{1}=-20^{\circ}$, respectively. As shown in Fig. 13, the interference signal is located in the side-lobe direction, leading to strong interference at the receiver. Fig. 14 presents the communication constellation diagram with interference suppression using the near-field model, compared to the counterpart scheme that only aligns with the target signal. It can be observed that the interference signal reduces the SINR and causes the constellation diagram to become more dispersed. Furthermore, the constellation diagram exhibits a pattern similar to that of non-orthogonal multiple access (NOMA), due to the power disparity between Tx. 1 and Tx. 2. However, the interference suppression scheme significantly improves the SINR, restoring the compactness of the constellation diagram. The theoretical and experimental beampatterns of both schemes are shown in Fig. 15. It is noteworthy that both the theoretical and experimental results presented in Fig. 15a and Fig. 15b demonstrate a consistent trend in interference suppression and target signal enhancement achieved through the 2-bit RISS. Specifically, a gain is observed in the direction of the target signal, while deep nulls are formed in the directions of interference. However, due to practical system imperfections such as phase shift deviations in the reflection elements and environmental disturbances, the actual level of interference suppression exhibits a noticeable deviation from the theoretical prediction. Additionally, it is also observed that the interference suppression scheme results in a slight reduction in the main-lobe gain compared to the only align scheme.
@@
(b) Robust interference suppression, $\Delta_{\text{EVM}}=6.5897\%$.
@@
Figure 15: The beampattern of align scheme and interference suppression scheme in the near-field model. The target signal is located in $\phi_{2}^{\text{azi}}=10^{\circ}$, while the interference signal is located in $\phi_{1}^{\text{azi}}=-20^{\circ}$.
In contrast, as illustrated in Fig. 16 and Fig. 17, the performance of target enhancement and interference suppression in the far-field model is weak. To better understand the underlying cause of this phenomenon, we present theoretical 2D heatmaps under both the near-field and far-field schemes. It can be observed from Fig. 18 that both schemes place the Tx.2 within the main-lobe while forming nulls in the directions of Tx.1. However, the near-field scheme exhibits fewer sidelobes and does not generate significant signal gain in the near region. In contrast, the far-field scheme produces more pronounced sidelobes and exhibits gain along the entire path at $\phi_{2}^{\text{azi}}=10^{\circ}$. In addition, there is a slight offset between the main-lobe direction and the actual position of the Tx.2. This misalignment is caused by the combined effects of the interference suppression constraint and the 2-bit RISS quantization. Furthermore, hardware imperfections, such as phase adjustment errors in the RISS, may also be contributing factors to the inferior performance of the far-field scheme compared to the near-field scheme.
@@
(b) Robust interference suppression, $\Delta_{\text{EVM}}=20.0157\%$.
@@
Figure 17: The beampattern of align scheme and interference suppression scheme in the far-field model. The target signal is located in $\phi_{2}^{\text{azi}}=10^{\circ}$, while the interference signal is located in $\phi_{1}^{\text{azi}}=-20^{\circ}$.
(a) Theoretical heatmap of near-field scheme.
(b) Theoretical heatmap of far-field scheme.
Figure 18: Heatmaps of both near-field and far-field schemes. The RISS is located at (0,0) and target signal is located in $\phi_{2}^{\text{azi}}=10^{\circ}$, while the interference signal is located in $\phi_{1}^{\text{azi}}=-20^{\circ}$.
Finally, to further investigate interference suppression, we deactivate Tx. 2, allowing the receiver to capture only the interference signal. The constellation measurements for both the align scheme and the interference suppression scheme using the near-field model are displayed in Fig. 19. With only Tx. 1 active, the constellation in Fig. 19a shows that the align scheme retains significant interference signal power. In contrast, the interference suppression scheme shows effective interference mitigation, resulting in a highly dispersed constellation diagram of the interference signal, further demonstrating the robustness of the interference suppression technique.
@@ VI Conclusion
Appendix A Convergence Analysis of the Proposed Algorithms
A-A Convergence analysis of the Proposed Algorithm 1
By observing Eq. (15), we note that $\bm{\mathcal{C}}_{1}$ and $\bm{\mathcal{C}}_{2}$ differ only by a constant term $P_{i},\forall i\in\{1,2\}$. For simplicity, we let $P_{1}=P_{2}$, and denote the common term as $\bm{\mathcal{C}}$, i.e., $\bm{\mathcal{C}}_{1}^{t}=\bm{\mathcal{C}}_{2}^{t}=\bm{\mathcal{C}}_{t}$. For the proposed IRM algorithm, we begin by formulating the Lagrangian function corresponding to problem (P1.2) as follows
$\mathcal{L}$ $=-\text{trace}(\bm{\mathcal{C}}_{t}\bm{\mathcal{A}}_{d})+\epsilon_{t}r+\bm{\mu}^{T}(\text{diag}(\bm{\mathcal{C}}_{t})-M\mathbf{1}_{N})$
$\qquad\qquad+\sum_{k\neq d}\lambda_{k}\left(\text{trace}(\bm{\mathcal{C}}_{t}\bm{\mathcal{A}}_{k})-\tau_{k}\right)$
$\qquad\qquad-\text{trace}\left(\bm{\Gamma}_{1}\left(r\mathbf{I}_{N-1}-\mathbf{V}_{t}^{H}\bm{\mathcal{C}}_{t}\mathbf{V}_{t}\right)\right)$
$\qquad\qquad-\text{trace}\left(\bm{\Gamma}_{2}\bm{\mathcal{C}}_{t}\right)+\lambda_{e}\left(\text{trace}(\bm{\mathcal{C}}_{t})-MN^{2}\right),$ (41)
where $\bm{\mu}\in\mathbb{R}^{N}$, $\lambda_{k}\geq 0$, $\bm{\Gamma}_{1}\in\mathbb{S}_{+}^{N-1}$, $\bm{\Gamma}_{2}\in\mathbb{S}_{+}^{N}$, and $\lambda_{e}\geq 0$ denote the dual multipliers associated with the equality constraints, inequality constraints, semidefinite constraints, and trace constraint, respectively. $\mathbb{R}^{N}$ denotes the space of $N$-dimensional real-valued vectors, and $\mathbb{S}_{+}^{N}$ denotes the cone of $N\times N$ real symmetric positive semidefinite matrices.
Rewriting the Lagrangian in a more compact form yields
$\mathcal{L}=$ $\text{trace}\left(\bm{\mathcal{C}}_{t}\bm{\Pi}\right)+r(\epsilon_{t}-\text{trace}(\bm{\Gamma}_{1}))-\bm{\mu}^{T}(M\mathbf{1}_{N})$
$\qquad\qquad\qquad\qquad-\sum_{k\neq d}\lambda_{k}\tau_{k}-\lambda_{e}MN^{2},$ (42)
where $\bm{\Pi}=-\bm{\mathcal{A}}_{d}+\text{diag}(\bm{\mu})+\sum_{k\neq d}\lambda_{k}\bm{\mathcal{A}}_{k}+\mathbf{V}_{t}\bm{\Gamma}_{1}^{H}\mathbf{V}_{t}^{H}-\bm{\Gamma}_{2}+\lambda_{e}\mathbf{I}_{N}$. The dual function is then defined as $g=\inf_{\bm{\mathcal{C}}_{t},r}\mathcal{L}$, which is finite if the stationarity conditions are satisfied. The first-order optimality conditions are given by
$\frac{\partial\mathcal{L}}{\partial r}=\epsilon_{t}-\text{trace}(\bm{\Gamma}_{1})=0,$
$\frac{\partial\mathcal{L}}{\partial\bm{\mathcal{C}}_{t}}=\bm{\Pi}=\mathbf{0}.$ (43)
Thus, the dual problem can thus be written as
$\min_{\bm{\mu},\bm{\lambda},\bm{\Gamma}_{1},\bm{\Gamma}_{2},\lambda_{e}}\quad$ $\bm{\mu}^{T}(M\mathbf{1}_{N})+\sum_{k\neq d}\lambda_{k}\tau_{k}+\lambda_{e}MN^{2}$ (44)
s.t. $\bm{\Pi}=\mathbf{0},$ (44a)
$\text{trace}(\bm{\Gamma}_{1})=\epsilon_{t},$ (44b)
$\lambda_{k}\geq 0,\bm{\Gamma}_{1}\succeq\mathbf{0},\bm{\Gamma}_{2}\succeq\mathbf{0},\lambda_{e}\geq 0.$ (44c)
Since $\bm{\mu}$ corresponds to equality constraints, it is not sign-restricted. By inspecting the primal problem, it can be verified that it is convex and satisfies Slater’s condition. Therefore, strong duality holds.
Let $(\bm{\mathcal{C}}_{t}^{*},r_{t}^{*})$ and $(\bm{\mathcal{C}}_{t+1}^{*},r_{t+1}^{*})$ denote the optimal solutions at iterations $t$ and $t+1$, respectively. The difference in the primal objective function values is given by
$p_{t+1}^{*}-p_{t}^{*}$ $=\left(-\text{trace}(\bm{\mathcal{C}}_{t+1}^{*}\bm{\mathcal{A}}_{d})+\epsilon_{t+1}r_{t+1}^{*}\right)$
$\qquad\qquad\qquad\qquad-\left(-\text{trace}(\bm{\mathcal{C}}_{t}^{*}\bm{\mathcal{A}}_{d})+\epsilon_{t}r_{t}^{*}\right)$
$=-\text{trace}\left((\bm{\mathcal{C}}_{t+1}^{*}-\bm{\mathcal{C}}_{t}^{*})\bm{\mathcal{A}}_{d}\right)+\epsilon_{t+1}r_{t+1}^{*}-\epsilon_{t}r_{t}^{*}.$ (45)
By substituting $\epsilon_{t+1}=\epsilon_{t}\varsigma$ and applying strong duality, i.e., $p_{t+1}^{*}-p_{t}^{*}=d_{t+1}^{*}-d_{t}^{*}$, we obtain
$\epsilon_{t}\varsigma r_{t+1}^{*}-\epsilon_{t}r_{t}^{*}=\underbrace{(d_{t+1}^{*}-d_{t}^{*})}_{\Delta_{d}}+\underbrace{\text{trace}\left((\bm{\mathcal{C}}_{t+1}^{*}-\bm{\mathcal{C}}_{t}^{*})\bm{\mathcal{A}}_{d}\right)}_{\Delta_{c}},$ (46)
where both $\Delta_{d}$ and $\Delta_{c}$ are bounded due to the compactness of the feasible set. Rearranging the expression leads to
$\varsigma r_{t+1}^{*}-r_{t}^{*}=\frac{\Delta_{d}+\Delta_{c}}{\epsilon_{t}}.$ (47)
Taking the limit as $t\to\infty$, and using the fact that $\epsilon_{t}=\epsilon_{0}\varsigma^{t}\to\infty$, we obtain
$\lim_{t\to\infty}\left(\varsigma r_{t+1}^{*}-r_{t}^{*}\right)=0.$ (48)
Since $r_{t}^{*}\to 0$ due to feasibility, the asymptotic behavior of the sequence satisfies
$r_{t+1}^{*}\approx\frac{r_{t}^{*}}{\varsigma}.$ (49)
Therefore, $r_{t}^{*}$ converges to zero at a linear rate of $\frac{1}{\varsigma}$.
Having established that $r_{t}^{*}$ converges to zero at a linear rate of $1/\varsigma$, we next show that the sequence $\bm{\mathcal{C}}_{t}^{*}$} converges to a locally optimal solution $\bm{\mathcal{C}}^{*}$, which satisfies $\text{rank}(\bm{\mathcal{C}}^{*})\leq 1$. Note that this conclusion aligns with the results in [32, Proposition 3.6].
Specifically, by the constraint $(c)$ and the fact that $\lim_{t\to\infty}r_{t}^{*}=0$, we have
$r_{t}^{*}\mathbf{I}_{N-1}-\mathbf{V}_{t}^{H}\bm{\mathcal{C}}_{t}^{*}\mathbf{V}_{t}\succeq\mathbf{0}\quad\Rightarrow\quad\lim_{t\to\infty}\mathbf{V}_{t}^{H}\bm{\mathcal{C}}_{t}^{*}\mathbf{V}_{t}=\mathbf{0}.$ (50)
[…]
$\bm{\mathcal{C}}_{t}^{*}=\alpha_{t}\mathbf{c}_{t}\mathbf{c}_{t}^{H},\quad\alpha_{t}\geq 0,\quad\|\mathbf{c}_{t}\|_{2}=1,$ (51)
where $\mathbf{c}_{t}$ is the normalized dominant eigenvector of $\bm{\mathcal{C}}_{t}^{*}$.
Now, consider the difference in the constraint projections between successive iterations
$\lim_{t\to\infty}\left(\mathbf{V}_{t+1}^{H}\bm{\mathcal{C}}_{t+1}^{*}\mathbf{V}_{t+1}-\mathbf{V}_{t}^{H}\bm{\mathcal{C}}_{t}^{*}\mathbf{V}_{t}\right)=\mathbf{0}.$ (52)
Since $\text{rank}(\bm{\mathcal{C}}_{t}^{*})=1$ asymptotically, we may assume
$\bm{\mathcal{C}}_{t+1}^{*}\approx\alpha\bm{\mathcal{C}}_{t}^{*},\quad\alpha\in\mathbb{R},$ (53)
which implies that $\bm{\mathcal{C}}_{t+1}^{*}$ and $\bm{\mathcal{C}}_{t}^{*}$ share the same eigenvectors. Hence, $\lim_{t\to\infty}\mathbf{c}_{t+1}=\mathbf{c}_{t}.$ As a consequence, the basis $\mathbf{V}_{t}$ becomes asymptotically constant, i.e., $\mathbf{V}_{t}\to\mathbf{V}^{*}$ as $t\to\infty$. Therefore, under the fixed $\mathbf{V}^{*}$, the subproblem (P1.2) converges to a locally optimal solution $\bm{\mathcal{C}}^{*}$ that satisfies the Karush-Kuhn-Tucker conditions (KKT) conditions.
A-B Convergence Analysis of the Proposed Algorithm 2
Note that the proposed AO algorithm can be regarded as a special case of the block coordinate descent (BCD) method. Specifically, the phase vector of the RISS, denoted as $\bm{\varphi}=\{\varphi_{1},\dots,\varphi_{N}\},\varphi_{k}\in[0,2\pi),\forall k\in N$, is naturally partitioned into $N$ scalar blocks. In each iteration of the AO algorithm, only one scalar variable $\varphi_{k}$ is optimized while keeping all other variables fixed, which aligns precisely with the coordinate-wise update rule in the BCD framework. Therefore, the proposed AO algorithm can be interpreted as a BCD method with scalar block updates and exact maximization along each coordinate. This observation enables us to leverage classical results from BCD theory to rigorously yet concisely support the convergence analysis.
The convergence of the AO algorithm is ensured by the following three properties
  • Compactness and boundedness: The domain of $\bm{\varphi}$ is compact, i.e., $\varphi_{k}\in[0,2\pi),\forall k\in N$, and the objective function $f(\bm{\varphi})=S_{2}-\eta S_{1}$ is continuous and upper-bounded over this domain. Specifically, $f(\bm{\varphi})\leq N$, ensuring the existence of a finite supremum.
  • Monotonic ascent: In each iteration, the objective function is exactly maximized over a single coordinate using a one-dimensional search. This guarantees that the objective value is non-decreasing across iterations, i.e., $f(\bm{\varphi}^{(t+1)})\geq f(\bm{\varphi}^{(t)})$.
  • Convergence to a stationary point: Since the sequence $f(\bm{\varphi}^{(t)})$} is monotonically non-decreasing and upper-bounded, it converges to a finite limit $f^{*}$. Moreover, any limit point $\bm{\varphi}^{*}=\{\varphi_{1}^{*},\dots,\varphi_{N}^{*}\}$ satisfies coordinate-wise optimality. That is, further improvement in the objective function is not possible by updating any individual coordinate $\varphi_{k}$. When the objective function is differentiable, this implies that the gradient vanishes along each coordinate direction, indicating that $\bm{\varphi}^{*}$ is a stationary point.
Based on these three conditions and standard results from BCD convergence theory[46], we conclude that the proposed algorithm converges to a stationary point. For a more detailed derivation, readers are referred to classical convergence proofs of BCD algorithms[46].
Appendix B Hardware Information Supplement
@@
Array Size $\times10$ units